OpenStudy (kriscoin):

The value of x in the proportion 1/2:2/3 = 3/4:x is 1 4/9 1 7/9 1/4

OpenStudy (texaschic101):

(1/2) / (2/3) = (3/4) / (x) cross multiply 1/2x = (2/3)(3/4) 1/2x = 1/2 x = 1 check.. (1/2) / (2/3) = (3/4) / (1/1) cross multiply (1/2)(1/1) = (2/3)(3/4) 1/2 = 1/2 (correct) so x = 1

OpenStudy (kriscoin):

Thank you!!!